ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 46, 57, 58 Kopylov, Yaroslav, Sobolev Institute of Mathematics, Novosibirsk, Russia Normal solvability of the operator of exterior derivation on some surfaces in ℝ^{N} The operator d of exterior derivation is studied on a smooth surface of revolution M in ℝ^{N}. The operator d is considered as an unbounded one defined on a subspace of the Banach space of measurable differential forms on M having their modulus integrable to the power p ≥ 1. Some conditions of normal solvability of d are presented (a linear operator is called normally solvable if its range is closed). Reference: Ya. A. Kopylov, On normal solvability of the operator of exterior derivation on a surface of revolution, Siberian Math. J. 38 (1997), No. 6, 1300–1307.